why does the inertial mass of an object change when the velocity of the object reaches the velocity of light .

Grade:10

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The mass at speeds close to the speed of light follows the rest mass concept.
m=\frac{m_{0}}{\sqrt{1-\frac{v^2}{c^2}}}
Hence forv\rightarrow c,m_0\rightarrow 0, so that the mass m is finite.
Regards
Sumit